# Deformable Mamba for Wide Field of View Segmentation

### v1

@@ 4.1 Datasets
To study the unification of wide-FoV segmentation, we conduct experiments on three 360° and two 180° datasets.
@@
SynPASS [67] is a virtual outdoor 360° dataset from CARLA [13]. It contains 9,080 images with a resolution of $1024{\times}2048$, covering various weather conditions (cloudy, foggy, rainy, sunny) and illumination (daytime, nighttime).
@@ 4.3 Quantitative Results
Given that previous works were largely constrained to specific Field-of-View (FoV) settings, which poses challenges for fair cross-dataset evaluation. Therefore, we selected representative state-of-the-art models as our baselines, including SegNeXt [21], which achieves superior performance on pinhole images, and 360SFUDA [73], which is specifically designed for panoramic scenarios. To ensure fair comparison, we carefully match the parameter counts across all implementations and conduct extensive evaluations on the aforementioned datasets with varying FoV characteristics.
@@
Results on 360° Stanford2D3D. Table 1 shows superior performance of our method compared to both panoramic [14, 51, 74, 66, 67, 53, 33, 73] and pinhole-based [21, 37] methods. Specifically, our model with the VMamba-T as backbone achieves a significant mIoU improvement of ${+}2.9\%$ to ${+}13.7\%$ over panoramic baselines.
Notably, compared to the VMamba-T + UperNet [57], our lightweighted DMamba-M with the same backbone achieves a ${+}2.5\%$ higher mIoU while reducing parameter count by 30% and FLOPs by 80%. Furthermore, our DMamba-T model, which has a comparable parameter count to VMamba-T + UperNet, demonstrates even better performance with a ${+}3.4\%$ mIoU improvement.
@@
Results on 360° Matterport3D.
While our DMamba-M demonstrates remarkable performance on Stanford2D3D dataset, the challenge becomes more pronounced on large-scale Matterport3D panoramic datasets, where we observe a relatively modest improvement margin of 0.3 mIoU over the previous SOTA model SegNeXt as shown in Table 1. Remarkably, our DMamba-T achieves more substantial gains of ${+}0.9\%$ and ${+}1.7\%$ mIoU compared to SegNeXt and VMamba-T respectively, with comparable model capacities.
Results on 360° SynPASS. Beyond real-world datasets, we evaluate the generalizability of our method on synthetic and outdoor scenes on the SynPASS dataset. It is worth noting that despite SegNeXt’s superior performance over 360SFUDA in real-world scenarios, it exhibits a substantial performance degradation on the synthetic dataset. This discrepancy underscores the sensitivity of certain architectures to distribution shifts. In contrast, our method, as shown in Table 2, achieves the state-of-the-art mIoU of 42.78%, surpassing all existing panoramic- and pinhole-based methods. Specifically, our DMamba-T overall achieves top scores on pedestrian, pole, traffic sign, static and dynamic. For traffic sign with slender properties, our model enjoys more than ${+}6.0\%$ mIoU.
@@
Results on 180° WoodScape. In Table 3, we present a comparison of our method with other models using 180° fisheye inputs from the WoodScape dataset [64]. Semi-supervised methods such as MeanTeacher [52], CPS [7], and FishSegSSL [43] show lower performance due to the limited data. Compared to multi-task models, our proposed DMamba, designed with deformable settings, outperforms all variants of the OmniDet [31] model.
Although SegNeXt and 360SFUDA perform well on the WoodScape dataset, our method, with a more compact architecture still achieves a higher mIoU by ${+}0.9\%$ and ${+}0.8\%$, respectively. The DMamba-T variant further advances the state-of-the-art mIoU to 83.9%, confirming DMamba’s effectiveness in capturing heavily distorted object information in wide-FoV segmentation.
@@
Results on 180° SynWoodScape. Due to the lack of baseline comparisons on SynWoodScape, we limited our evaluation to SegNeXt and 360SFUDA. Similar to the results on SynPASS, SegNeXt and 360SFUDA perform similarly on real 180° datasets, while the former exhibits a significant performance degradation on synthetic 180° data. This further confirms the challenge of achieving consistently high performance across diverse datasets. In comparison, our DMamba-M of similar scale outperforms them by ${+}4.36\%$ and ${+}1.73\%$ mIoU, respectively, achieving an improved segmentation accuracy of 69.20 mIoU on SynWoodScape.
@@
4.4 Qualitative Results
To demonstrate the improved performance of our model over Mamba-based models, we compared the visualization results of DMamba and VMamba on the Stanford2D3D and SynWoodScape. As discussed in Sec. 3.1 and  3.4, the scanning approach in Mamba-based models introduces biases in visual tasks that limit their effectiveness in dense prediction tasks. These biases increase randomness and reduce the model’s ability to capture local spatial relationships, which is particularly evident in the recognition and segmentation of smaller objects. As shown in Fig. 5, the baseline model struggles to accurately segment narrow and elongated areas such as door frames, vegetation contours, traffic light poles, and lane markings. In contrast, our proposed DMamba effectively addresses these challenges, significantly improving segmentation performance on small and elongated objects.
Table 5: Ablation study for analysis of backbones (BB) and decoders (Dec) on the 360° dataset Stanford2D3D.
[…]
BB
Dec
Total
BB
Dec
Total
mIoU
mAcc
ResNet-UperNet ResNet50 [23] UperHead [57] 44.3 23.5 40.5 64.0 22.9 250.7 273.6 53.34 62.42
ResNet-UperFormer ResNet50 [23] MusterHead [60] 14.3 23.5 203.1 226.6 22.9 211.5 234.4 54.39 63.27
ResNet-CGRSeg ResNet50 [23] CGRHead [41] 58.0 23.5 282.6 306.1 22.9 31.4 54.3 54.12 62.94
ResNet-DMamba (Ours) ResNet50 [23] DMambaHead 27.9 23.5 77.3 100.8 22.9 38.8 61.7 57.22 66.32
Swin-UperNet Swin-T [38] UperHead [57] 34.9 27.5 31.5 59.0 25.6 211.5 237.1 53.82 63.06
Swin-UperFormer Swin-T [38] MusterHead [60] 30.2 27.5 19.1 46.6 25.6 21.9 47.5 52.36 54.54
Swin-CGRSeg Swin-T [38] CGRHead [41] 52.2 27.5 40.6 68.1 25.6 5.5 31.1 54.26 62.90
Swin-DMamba (Ours) Swin-T [38] DMambaHead 37.3 27.5 11.2 38.7 25.6 6.5 32.1 55.09 65.30
VMamba-UperNet VMamba-T [37] UperHead [57] 35.0 29.5 31.5 61.0 25.3 206.9 232.2 56.80 65.01
VMamba-UperFormer VMamba-T [37] MusterHead [60] 29.8 29.5 19.1 48.6 25.3 21.4 46.7 58.12 66.97
VMamba-CGRSeg VMamba-T [37] CGRHead [41] 42.2 29.5 40.5 70.0 25.3 5.0 30.3 56.45 65.07
DMamba-M (Ours) VMamba-T [37] DMambaHead 40.0 29.5 11.1 40.6 25.3 6.0 31.3 59.30 68.99
DMamba-T (Ours) VMamba-S [37] DMambaHead 17.9 49.4 11.1 60.5 44.7 6.0 50.7 60.18 71.68
@@ 4.5 Ablation Study
Effect of Backbone. In Table 5, we ablate three different kinds of commonly-used model backbones, namely CNN-based ResNet [23], ViT-based Swin Transformer [38] and Mamba-based VMamba [37]. For the sake of a fair comparison, we choose the backbones with a comparable number of learnable parameters and FLOPs. According to Table 5 it can be observed that the models equipped with VMamba [37] outperform other models consistently while CNN-based ResNet50 and ViT-based Swin Transformer achieve similar performance. This conclusion can be validated by employing the same decoder, such as UperHead [57], MusterHead [60], CGRHead [41], and our proposed DMambaHead, along with a consistent model backbone to perform the segmentation task. Notably, the combination of the VMamba [37] backbone with DMambaHead, resulting in our Deformable Mamba model, demonstrates superior performance compared to other variant combinations by a maximum of over $+6.9\%$ mIoU.
Effect of Decoder. Apart from the backbones, we also ablate the decoders. For every backbone, we conduct four experiments with different decoders, i.e., UperHead [57], MusterHead [60], CGRHead [41], and our proposed DMambaHead. All results are reported in Table 5. Using the same model backbone, our proposed DMambaHead outperforms other widely-used decoders by a $+3\%$ mIoU while a complexity degradation of $+20\%$, making the Deformable Mamba and its variants extremely practical in real-time wide-FoV accurate segmentation. It’s worth noting that the DMambaHead with VMamba-T [37] backbone achieves a drop of FLOPs from $232.2$G to $31.3$G, a drop of the learnable parameters from $61$M to $40.6$M while the performance of Deformable Mamba increases from a $+56.8\%$ to $+59.3\%$ mIoU compared with the VMamba-T [37] and UperHead [57] combination. The reason behind this is that our proposed DMambaHead is specifically designed for wide-FoV segmentation. Our deformable design enables the model to focus on the distortion introduced by the wide-FoV fisheye or panoramic images. This ablation proves that not only the VMamba [37] backbone but also the proposed DMambaHead contributes to the superior performance and efficiency of the Deformable Mamba model.
(a) Results on Matterport3D.
(b) Results on Stanford2D3D.
Figure 6: Analysis of segmentation with different wide FoVs.
Analysis of FoVs. To further analyze the wide-FoV segmentation, we conduct omni-directional analysis. As shown in Fig. 6, there are six different FoVs, ranging from 60° to 360°. Compared to the baseline VMamba [37], our DMamba methods obtain relative better results on all different FoVs. Note that the 60° performance of all methods are lower than other degrees on the Stanford2D3D dataset. One reason is that the object distribution and room layout in the 60° direction are different between Stanford2D3D and the Matterport3D datasets. Nonetheless, our Mini- and Tiny-version methods show consistent improvements over the baseline on both datasets in different FoVs.
@@ 5 Conclusion
In this work, we are the first to explore the unification of the wide field of view (FoV) segmentation tasks, including 180° and 360° image segmentation. Most existing models are tailored for pinhole camera images, making it challenging to extend their success to wide-FoV domains due to the severe image distortion and object deformation. Unlike previous approaches that require a specific architecture for each FoV or dataset, we propose a unified framework, Deformable Mamba, built on the emerging State Space Model (SSM) mechanism. Our framework is constructed with a flexible deformable decoder, which can be easily adapted to various models and datasets. Consistent improvements across five diverse wide-FoV datasets, ranging from indoor to outdoor, from synthetic to real-world scenes, demonstrate that our unified framework effectively addresses the unique challenges of wide-FoV segmentation.
Limitations and Future Work. Using large language models (LLMs) for unifying the wide-FoV segmentation is a potential solution but not discussed in this work, while LLMs can further include the segmentation methods that use vision-language architectures. Besides, cross-domain evaluation, e.g., transferring from a synthetic to a real wide-FoV scene, or from a narrow-FoV camera to a wide-FoV camera, will be further explored in our future work.

### v2

@@ 4 Experiment Results
To validate our deformable Mamba method, we compare with state-of-the-art approaches on three 360° panoramic and two 180° fisheye datasets. Additionally, we conduct experiments using CNN-, Transformer-, and Mamba-based architectures (e.g., ResNet, Swin, VMamba) to validate the efficiency and effectiveness of our Deformable Mamba Decoder.
@@ 4.1 Datasets
SynPASS [65] is a synthetic outdoor 360° dataset from CARLA [14]. It contains 9,080 images with a resolution of $1024{\times}2048$, covering various weather conditions (cloudy, foggy, rainy, sunny) and illumination (daytime, nighttime).
@@ 4.2 Implementation Details
4.3 Effectiveness and Robustness analysis
Table 1: Results on Stanford2D3D, using ResNet-50 [22] as the backbone with 23.5M parameters, 22.9G FLOPs and hierarchical embedding dimensions of {256, 512, 1024, 2048}.
[…]
mIoU Acc
UperHead [54] (CVPR20) 40.5 250.7 53.34 62.42
MusterHead [57] (CVPR21) 203.1 211.5 54.39 63.27
CGRHead [38] (CVPR23) 282.6 31.4 54.12 62.94
Ours 77.3 38.8 57.22 66.32
Table 2: Results on Stanford2D3D, using Swin-T [35] as the backbone with 27.5M parameters, 25.6G FLOPs and hierarchical embedding dimensions of {96, 192, 384, 768}.
[…]
mIoU Acc
UperHead [54] (CVPR20) 31.5 206.9 53.82 63.06
MusterHead [57] (CVPR21) 19.1 21.4 52.36 54.54
CGRHead [38] (CVPR23) 40.6 5.0 54.26 62.90
Ours 11.2 6.0 55.09 65.30
Table 3: Results on Stanford2D3D, using VMamba-T [34] as the backbone with 29.5M parameters, 25.3G FLOPs and hierarchical embedding dimensions of {96, 192, 384, 768}.
[…]
mIoU Acc
UperHead [54] (CVPR20) 31.5 206.9 56.80 65.01
MusterHead [57] (CVPR21) 19.1 21.4 58.12 66.97
CGRHead [38] (CVPR23) 40.6 5.0 56.45 65.07
Ours 11.2 6.0 59.30 68.99
Adaptability to different backbones. As mentioned in Sec. 1 and Sec. 3, decoupling the deformable setting from the overall framework and introducing it exclusively from the decoder perspective somewhat reduces the model’s capability to perceive distortions in wide-FoV data. The advantage, however, lies in enhancing the model’s flexibility. Through the plug-in distortion-aware decoder, we enable the model to adaptively perceive image deformations for different backbones. On the Stanford2D3D panoramic dataset, we compared our Deformable Mamaba Decoder with UperHead [54], MusterHead [57], CGRHead [38], evaluating their adaptability to CNN-, Transformer-, and Mamba-based backbones. As shown in Tables 1, 2, and 3, compared to decoders lacking distortion perception, our method achieves the best performance across ResNet-50 [22], Swin-T [35], and VMamba-T [34] backbones, which demonstrates the significance of distortion perception for wide-FoV image understanding.
Efficiency of the decoder. We emphasize that an efficient decoder is crucial for alleviating the overall model complexity. For instance, Liu et al. ultilize UperHead as the decoder to evaluate VMamba’s application in semantic segmentation tasks. As shown in Table 3, with VMamba as the backbone, the model has only 25.3G FLOPs, while the decoder UperHead, with 206.9G FLOPs, becomes the bottleneck of the models’ computational complexity. While MusterHead and CGRHead reduce computational complexity, they exhibit strong sensitivity to the model scale. As evidenced in Table 1, their parameter count experiences substantial growth with increasing embedding dimensions. In contrast, our proposed decoder showcases superior compatibility with diverse model scales while retaining lower temporal and spatial complexity.
Based on the results, our method provides substantial distortion-aware capability for CNN-, Transformer-, and Mamba-based models while maintaining excellent efficiency, demonstrating its adaptability and robustness.
4.4 Qualitative Results
Visualization of distortion-aware capability. To showcase the distortion-aware capability of our decoder, we present a visual comparison against UperHead [54] lacking deformable desin on the 360° Stanford2D3D and 180° SynWoodScape datasets. As shown in Fig. 5, the baseline model struggles to accurately segment narrow and elongated areas such as door frames, vegetation contours, traffic light poles, and lane markings. In contrast, our proposed DMamba with Deformable Mamba Decoder effectively addresses these challenges, significantly improving segmentation performance on small and elongated objects.
Analysis of feature maps. As shown in Figure 6, to illustrate how the DMF module enhances distortion awareness, we visualize the channel map extracted from the encoder and the corresponding output features of DMF block. Specifically, we present the results of channels #$\#6$ and #$\#68$ in the second layer of the first decoder stage. Prior to passing through the DMF block, the features primarily focus on global structures while lacking adequate emphasis on crucial regions such as small objects in cluttered indoor environments 6(a) or severely distorted areas in wide-FoV outdoor scenes 6(b). Without explicit distortion-aware adaptation, the feature representation remains less effective in preserving fine-grained spatial details and compensating for geometric distortions. The model shows enhanced attention to objects at different scales after traversing the DMF block.
(a) Visualization of feature maps in indoor scenarios
(b) Visualization of feature maps in outdoor scenarios
Figure 6: Visualization of feature maps. From left to right: features extracted by Mamba-based encoder (Before DMF) and features fused by the DMF module (After DMF). Effective detection and recognition of small objects in indoor scenarios (a), such as pictures on walls, requires the model to exhibit enhanced local perception capabilities. In contrast, outdoor scenarios (b) captured by wide-FoV cameras are often characterized by significant object deformation and distortion. To facilitate accurate scene understanding in such cases, the model must be equipped with a robust ability to handle and perceive distortion effectively.
@@ 4.5 Quantitative Results
We extensively evaluate our proposed method across the five aforementioned datasets, addressing the critical challenge of data heterogeneity prevalent in existing methods. Through meticulous comparative experiments with state-of-the-art approaches and a careful re-implementation of [20, 70], we systematically investigate the performance variations induced by diverse data distributions and types, thereby providing comprehensive insights into model generalizability and robustness.
@@
Results on 360° Stanford2D3D dataset. Table 4 shows superior performance of our method compared to both distortion-aware [15, 48, 71, 64, 65, 50, 31, 70] and non-distortion-aware [20, 34] methods. Specifically, our DMamba-M model with the VMamba-T as backbone achieves a significant mIoU improvement of ${+}2.9\%$ to ${+}13.7\%$ over panoramic baselines. Notably, compared to the VMamba [54], our lightweighted decoder with the same backbone achieves a ${+}2.5\%$ higher mIoU while reducing overall parameters by 20% and FLOPs by 86%. Furthermore, our DMamba-T model, which has a comparable parameter count to VMamba, demonstrates even better performance with a ${+}3.4\%$ mIoU improvement.
Results on 360° Matterport3D dataset. While our DMamba-M demonstrates remarkable performance on Stanford2D3D dataset, the challenge becomes more pronounced on large-scale Matterport3D panoramic datasets, where we observe a relatively modest improvement margin of ${+}0.3\%$ mIoU over the previous SOTA model SegNeXt as shown in Table 4. Remarkably, our DMamba-T achieves more substantial gains of ${+}0.9\%$ and ${+}1.7\%$ mIoU compared to SegNeXt and VMamba-T respectively, with comparable model capacities.
@@
Results on 360° SynPASS dataset. Beyond real-world datasets, we evaluate the generalizability of our method on synthetic and outdoor scenes on the 360° SynPASS dataset. It is worth noting that despite SegNeXt’s superior performance over 360SFUDA in real-world scenarios, it exhibits a substantial performance degradation on the synthetic dataset. This discrepancy underscores the sensitivity of certain architectures to data distribution shifts. In contrast, our method, as shown in Table 5, achieves the state-of-the-art mIoU of 42.78%, surpassing all existing panoramic- and pinhole-based methods. Explicitly, our DMamba-T overall achieves top scores on pedestrian, pole, traffic sign, static and dynamic. For traffic sign with slender properties, our model enjoys more than ${+}6.0\%$ mIoU.
Results on 180° WoodScape dataset. In Table 6, we present a comparison of our method with other models using 180° fisheye inputs from the WoodScape dataset [62]. Semi-supervised methods such as MeanTeacher [49], CPS [8], and FishSegSSL [40] show lower performance due to the limited data. Compared to multi-task models, our proposed DMamba, designed with deformable settings, outperforms all variants of the OmniDet [29] model. Although SegNeXt and 360SFUDA perform well on the WoodScape dataset, our method, with a more compact architecture still achieves a higher mIoU by ${+}0.9\%$ and ${+}0.8\%$, respectively. The DMamba-T variant further advances the state-of-the-art mIoU to 83.9%, confirming DMamba’s effectiveness in capturing heavily distorted object information in wide-FoV segmentation.
Results on 180° SynWoodScape dataset. Due to the lack of baseline comparisons on SynWoodScape, we limited our evaluation to SegNeXt and 360SFUDA. As shown in Table 7, similar to the results on SynPASS, SegNeXt and 360SFUDA perform similarly on real 180° datasets, while the former exhibits a significant performance degradation on synthetic data. This further confirms the challenge of achieving consistently high performance when data distribution shifts. In comparison, our DMamba-M of similar scale outperforms them by ${+}4.36\%$ and ${+}1.73\%$ mIoU, respectively, achieving an improved segmentation accuracy of 69.20 mIoU on SynWoodScape.
@@ 4.6 Ablation Study
To further analyze the effect of different components on our proposed framework, we conduct ablation study on the Stanford2D3D dataset.
Analysis of scanning mechanism. To ensure consistency, we adopt the same scanning method as in [34], as our model primarily uses its Mamba-based encoder. As mentioned in Sec. 1, the SSM-based Mamba processes input tokens sequentially, which is suitable for textual data. However, for image data, where tokens have spatial relationships, multi-directional scanning is more appropriate. To highlight the difference between various scanning methods on our framework, we conducted ablation experiments on the Stanford2D3D dataset. As shown in Table 8, both uni-directional scanning and bi-directional scanning results in slight performance drop.
Table 8: Analysis of scanning methods.
        Scan Method       mIoU       mAcc
      Uni-direction       57.53       67.52
      Bi-direction       57.94       67.66
      Quadri-direction (Ours)       59.30       68.99
Study of defomable convolution. In Sec. 1 and Sec. 2, we outlined the limitations of tightly coupling deformable designs with the entire model and the challenges faced by Mamba-based models in integrating deformable mechanisms due to their sequential processing of input tokens. We propose to introduce sufficient distortion awareness into the model by paralleling DCN in the decoder and integrating it through feature fusion. We replace deformable convolution with traditional one as shown in Table 9, the results prove the effectiveness of this trade-off solution.
Table 9: Ablation Study of deformable designs.
  Deformable mIoU mAcc
57.45 66.52
59.30 ($+1.85\%$) 68.99 ($+2.47\%$)
Analysis of upsampling methods. Sampling methods play a critical role in our approach. As mentioned earlier, our decoder utilizes two branches to achieve complete feature modeling and the incorporation of distortion awareness respectively. To optimize computational efficiency, we adopt a single-layer module in each branch, which reduces model complexity while limiting the model’s learnable capacity. As a result, we disregard parameter-free interpolation-based upsampling methods, opting instead to fuse features from both branches and employ periodic arrangement to preserve all information. In Table 10, we compare bilinear and bicubic methods, showing the importance of our design.
Table 10: Analysis of upsample methods.
        Upsample Method       mIoU       mAcc
      Bilinear Interpolation       56.90       66.08
      Bicubic Interpolation       56.97       66.12
      PixelShuffle (Ours)       59.30       68.99
@@ 5 Conclusion
In this work, we are the first to explore the unification of the wide field of view (FoV) segmentation tasks, including 180° and 360° image segmentation. Most existing models are tailored for pinhole camera images, making it challenging to extend their success to wide-FoV domains due to the severe image distortion and object deformation. We incorporate the model’s distortion-aware capability exclusively from the decoder perspective, decoupling it from the overall architecture to enhance flexibility and facilitate seamless integration with various model architectures. Consistent improvements across five diverse wide-FoV datasets, ranging from indoor to outdoor, from synthetic to real-world scenes, demonstrate that our method effectively addresses the unique challenges of wide-FoV segmentation.
Limitations and Future Work. Given the rapid advancement of Large Language Models(LLMs), integrating multi-modal LLMs presents a promising direction for future research. Their strong generalization capabilities and semantic understanding could significantly enhance our unified semantic segmentation framework, particularly in handling diverse wide-FoV segmentation tasks. This integration could bridge the gap between linguistic semantic understanding and geometric feature learning, potentially leading to more robust and versatile segmentation models.